OpenStudy (anonymous):

what is 18/30 in lowest terms a. 3/5 b. 4/15 c. 3/10 d. 2/5

OpenStudy (anonymous):

@texaschic101 @TeenZombie46 @KatnissStyles000 @Christian_10_ @ElisaNeedsHelp please help

OpenStudy (anonymous):

3/5

OpenStudy (anonymous):

18 / 30 = (2 * 32) / (2 * 3 * 5) = ((2 * 32) : (2 * 3)) / ((2 * 3 * 5) : (2 * 3)) = (18 : 6) / (30 : 6) = 3 / 5

OpenStudy (anonymous):

ok um thank u i will fan and medal but there is 7 more :/

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

witch fraction is equivalent to 6/9? a.3/4 b.18/36 c.12/18 d. 1/3

OpenStudy (anonymous):

c

OpenStudy (anonymous):

what value of A makes the equation true ? 24/40=9/A A. 15 B. 12 C. 6 D. 25

OpenStudy (anonymous):

12

OpenStudy (anonymous):

no 1524/ 40 = 0.6 9 / 15 = 0.6

OpenStudy (anonymous):

15*

OpenStudy (anonymous):

3/8 x 5/6 = a.9/20 b. 5/16 c.4/7 d.1/6

OpenStudy (anonymous):

24/ 40 = 0.6 9 / 15 = 0.6

OpenStudy (anonymous):

o sorry

OpenStudy (anonymous):

?

OpenStudy (anonymous):

0.3125 = 3125/10000 = 5/16

OpenStudy (anonymous):

what is the reciprocal of 7/12? -7/12 7/12 12/7 -12/7

OpenStudy (anonymous):

12/7

OpenStudy (anonymous):

what is 1/4 divided by 3/5?

OpenStudy (anonymous):

It is the same thing as 3/5 multiplied by 4/1. 12/5, or 2 2/5

OpenStudy (anonymous):

what is the LCD of 1/6 and 3/8? 24 48 6 8

OpenStudy (anonymous):

The common denominator is 24: 1/6 = 4/24 and 3/8 = 9/24

OpenStudy (anonymous):

so 24?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thank u so much i will tell u what i get

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

glad to help

OpenStudy (anonymous):

;-; failed...

OpenStudy (anonymous):

WHAT

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

but how

OpenStudy (anonymous):

is that k12

OpenStudy (anonymous):

take the medal back

OpenStudy (anonymous):

missed 3

OpenStudy (anonymous):

what was that

OpenStudy (anonymous):

yes its k12 and no i said i would give u medal so i did

OpenStudy (anonymous):

no really i dont deserve it

OpenStudy (anonymous):

65%

OpenStudy (anonymous):

yes u do

OpenStudy (anonymous):

it did that to me

OpenStudy (anonymous):

lots of time

OpenStudy (anonymous):

u do k12?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

what grade?

OpenStudy (anonymous):

6th

OpenStudy (anonymous):

i know about this stuf

OpenStudy (anonymous):

im 7th and have no idea

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

when i do vocabulary checkpoint i get 75% and when i check what i did wrong it shows i didnt put the answer for those question

OpenStudy (anonymous):

i think thats what happened :/

OpenStudy (anonymous):

yea hate it i think its a glitch

OpenStudy (anonymous):

yea :/

OpenStudy (anonymous):

do u have twitter pic collage skype..?

OpenStudy (anonymous):

i have twitter

OpenStudy (anonymous):

i feel bad for you im sure im gonna fail my checkpoint to on irregular shapes area

OpenStudy (anonymous):

i hate math not my thing

OpenStudy (katnissstyles000):

um, sorry

OpenStudy (anonymous):

?